Find the convolution of the following functions. After integrating find the LaPlace transform of the convolution

f(t)=t^2

g(t)=e^-t
Answer #1

t. o f(t) = t get) et (f*9) (4)= 1800 gt-1)d7= 2 (4-g?d7 = 1² 24-22+ + 2 2f8g) = MAGYARIHUANETA Jest (t2t-22 +2) dt = 0 -33-5